Rashed plans to take a loan of 100,000 AED to buy a car, the bank available loan options are:   
                                                                                                            
Option 1: Total amount owed is due with a single payment at the end of year 10 with 7.5 % simple interest per year.
Option 2: Total amount owed is due with a single payment at end of year 8 with 6.5% annually compound interest.

Which option should Rashed choose? Match the closest correct answers for the below questions:

      -       A.       B.       C.       D.       E.       F.       G.   

The total amount owed using Option 1?

      -       A.       B.       C.       D.       E.       F.       G.   

The total amount owed using Option 2?

      -       A.       B.       C.       D.       E.       F.       G.   

Which account Rashed should select?

A.

[20,000]

B.

[Option 1]

C.

[165,499]

D.

[60,000]

E.

[160,000]

F.

[Option 2]

G.

[175,000]
Answer #1

(1) (G)

Total amount owed using option 1 (AED) = Loan value + (Loan value x Number of years x Annual Interest rate)

= 100,000 + (100,000 x 10 x 7.5%) = 100,000 + 75,000 = 175,000

(2) (C)

Total amount owed using option 2 (AED) = Loan value x (1 + Annual Interest rate)Number of years

= 100,000 x (1.065)8 = 100,000 x 1.65499 = 165,499

(3) (F)

Rasheed should choose Option 2 because this has the value of total amount owed is lower.